# A Bag of Tricks for Few-Shot Class-Incremental Learning

### v1

@@ 5 Conclusion
iCaRL [26] 68.68 52.65 48.61 44.16 36.62 29.52 27.83 26.26 24.01 23.89 21.16
EEIL [2] 68.68 53.63 47.91 44.20 36.30 27.46 25.93 24.70 23.95 24.13 22.11
TOPIC [34] 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28
Rebalancing [13] 68.68 57.12 44.21 28.78 26.71 25.66 24.62 21.52 20.12 20.06 19.87
SPPR [43] 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33
MetaFSCIL[8] 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64
F2M [31] 81.07 78.16 75.57 72.89 70.86 68.17 67.01 65.26 63.36 61.76 60.26
CEC [37] 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28
FACT [40] 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94
LIMIT [41] 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41
SoftNet [14] 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75
ALICE [25] 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10
SAVC [33] 81.85 77.92 74.95 70.21 69.96 67.02 66.16 65.30 63.84 63.15 62.50
Ours 82.31 78.03 75.45 70.99 71.06 67.85 67.44 66.05 64.95 64.31 63.75
@@
iCaRL [26] 71.77 61.85 58.12 54.60 51.49 48.47 45.90 44.19 42.71
Rebalancing [13] 72.30 66.37 61.00 56.93 53.31 49.93 46.47 44.13 42.19
TOPIC [34] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
EEIL [2] 61.31 46.58 44.00 37.29 33.14 27.12 24.10 21.57 19.58
FSLL [22] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS [22] 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M [31] 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84
CEC [37] 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63
MetaFSCIL [8] 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19
C-FSCIL [12] 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41
FACT [40] 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49
CLOM [44] 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00
LIMIT [41] 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19
SoftNet [14] 79.77 75.08 70.59 66.93 64.00 61.00 57.81 55.81 54.68
ALICE [25] 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70
SAVC [33] 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11
Ours 84.3 79.59 75.49 71.4 68.45 65.129 62.2 60.52 59.57
S3 Comparison to State-of-the-art
In this section, we present the results for CUB-200 and miniImageNet datasets and their comparison to prior works. As we observe from Table S1, our framework outperforms all existing works on CUB-200 and sets a state-of-the-art of 63.75%. Similarly, for miniImageNet in Table S2, our framework performs better than all prior works, showing a new best accuracy of 59.57%.
S4 Additional Results
S4.1 Baseline
@@
Table S7: Performance of prior works and our framework for different encoder sizes.
Method CIFAR-100 CUB-200 MiniIN
SAVC [33] w/ $*{}^{*}$ 53.12 62.50 57.11
SAVC [33] w/ ResNet-50 49.30 56.92 56.92
SoftNet [14] w/ ResNet-18 55.33 56.75 54.68
SoftNet [14] w/ ResNet-50 55.65 56.15 54.25
Ours w/ ResNet-18 58.55 63.60 59.11
Ours w/ ResNet-50 58.75 63.75 59.57
*SAVC employed ResNet-20 for CIFAR-100 and ResNet-18 for CUB-200 and
miniImageNet.
Table S8: Comparison of computational complexity with a batch-size of 64.
Method Acc. Train (base) Train (inc.) Inference
SAVC 53.12 810 950 2150
Ours (ResNet-18) 58.55 630 935 2150
Ours (ResNet-50) 58.75 450 710 1800
@@
S4.4 Training Tricks
Finally, we discuss the experiments on the training tricks. In Table S6, we show the results for training different encoders on all the datasets. As we observe in this table, increasing the model size in our framework improves the performance up to a certain model size. Notably, we observe that the best performance for all the datasets is obtained with ResNet-50. Since prior works SAVC [33] and SoftNet [14] originally reported their final performances with ResNet-18 or ResNet-20 backbones, we also report their performance using ResNet-50 for a fair comparison to our approach. For this experiment, we follow the exact implementation details reported in the original papers [33, 14]. As seen in Table S7, SAVC with ResNet-50 exhibits significant drops in performance on all datasets. SoftNet performs slightly better with ResNet-50 on CIFAR-100 but shows drops in performance for the other two datasets. Overall, the performance of these methods is significantly lower than that of our proposed framework when using the same encoder. We also report the performance of our framework using the ResNet-18 backbone, demonstrating considerably better results compared to other methods, irrespective of the encoder size.
S4.5 Performance on Different Shots
@@
SAVC [33] 79.85 71.29 64.11 57.08 53.45 50.45 48.74 46.54 44.45
SoftNet [14] 79.88 72.01 65.01 59.01 55.01 52.01 50.01 48.01 46.01
@@
SAVC [33] 79.85 72.00 65.45 59.17 55.25 52.2 50.16 48.11 46.94
SoftNet [14] 79.88 73.02 66.01 60.01 56.01 53.01 51.01 49.01 47.01
@@
SAVC [33] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
SoftNet [14] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
@@
SAVC [33] 79.85 75.45 72.18 68.26 65.66 62.15 60.27 58.19 56.01
SoftNet [14] 79.88 77.28 73.15 69.34 66.14 63.38 61.27 59.33 57.14
@@
S5 Time Complexity
In Table S8, we discuss the time complexity of our framework using a single Nvidia RTX 2080 GPU in comparison to SAVC [33]. To this end, we report throughput in frames per second (FPS), which is a common metric for indicating computational complexity. During the training phase on the base session (with the ResNet-50 backbone), the throughputs for SAVC and our framework are 810 FPS and 450 FPS, respectively. In this phase, our framework is relatively slower than SAVC. However, we observe a 1.4 times faster training speed when we use the ResNet-18 backbone in our framework, which comes at a 0.2% drop in performance. Nonetheless, during incremental training and inference, our framework with ResNet-18 is as fast as SAVC. Consequently, once trained, our framework can be reliably deployed in real-world applications with the same inference time as previous state-of-the-art while achieving enhanced performance.

### v2

@@ 4.2 Evaluation Protocols
4.3 Main Results
In this section, we present the main results of our tricks by cumulatively adding them to the baseline. The results are presented in Table 1, which include the category of the tricks, along with the stages at which the tricks are applied and the performance on CIFAR-100, CUB-200 and miniIN. As outlined in Section 3.2, we adopt the incremental-frozen framework as the baseline for our study. As shown in Table 1, the accuracy of this baseline for CIFAR-100, CUB-200, and miniImageNet datasets are 43.77%, 59.88%, and 45.08%, respectively. To ensure the best performance for the baseline, we perform an extensive hyper-parameter study presented in Appendix S3.1. As we later compare our results with prior works (Section 4.5), we observe that this baseline outperforms many of the prior studies, showing the robustness of our baseline.
Table 1: Impact of different tricks on the baseline for all datasets. Here, the tricks are added cumulatively to the baseline, and the ‘Stage’ indicates which stage of training the tricks are applied.
Category Trick Stage CIFAR-100 CUB-200 miniIN
-- Baseline -- 43.77 59.88 45.08
Stability + SupCon Base+Incremental 50.16 60.38 48.90
+ ETF vector Base 51.10 60.74 49.73
       + Pseudo-classes Base 51.65 62.18 54.79
Adaptability           + Inc. SubNet tuning Incremental 58.12 63.10 57.85
Training              + Pre-training Pre-training 58.34 63.30 58.01
                + Additional signal Base 58.55 63.60 59.11
Stability Tricks. Next, we discuss the results of including stability tricks into the baseline, beginning with the addition of SupCon loss. The results from this study show substantial improvement for all datasets (Table 1), obtaining accuracies of 50.16%, 60.38%, and 48.90%, for CIFAR-100, CUB-200, and miniImageNet respectively. As discussed earlier, a key element for ensuring high stability in FSCIL is to ensure increased separability in the embedding space (increased inter-class distance and reduced intra-class dispersion). In Figures 2(a) and 2(b), we plot the cumulative probability of inter-class and intra-class distances (defined in Section 4.2) for different tricks. We plot the cumulative probability of inter-class and intra-class distance (Song et al., 2023) instead of the average since the average distance can be misleading due to the presence of outliers, whereas the cumulative probability provides a more robust and nuanced representation of the distribution of distances. As we observe from Figure 2(a), adding SupCon provides a large increase in the inter-class distance compared to training with cross-entropy loss in the baseline. At the same time, SupCon greatly reduces the intra-class distance compared to the cross-entropy-based baseline (illustrated in Figure 2(b)). However, our findings diverge from those of Song et al. (2023), which indicated that while supervised contrastive loss effectively reduces intra-class distances, it unexpectedly leads to a reduction in inter-class distances as well. The divergence in our findings may be attributed to implementation specifics. For example, SAVC utilized multi-crop augmentation, which we do not use. Also, we adopted a different set of hyper-parameters determined through our study (Appendix S3.2).
We also show the overall class separation degree for different tricks in Figure 2(c). Class separation degree refers to the degree of distinctiveness or separability between different categories within the embedding space, which can be measured as defined in Section 4.2. It typically ranges between 0 and 1, with higher values indicating clearer boundaries and better separability between classes. As we observe, adding SupCon provides better class separation for both base and novel classes. Overall, the higher separation in the embedding space contributes to the large improvement in the overall accuracy. Analyzing the performance of both base and novel classes in Figure 2(d), we find improvements for both, with a 5.4% boost in base class accuracy and a 7.8% increase in novel class accuracy.
(a) Baseline.
(b) bag of tricks.
Our adaptability trick with Incremental SubNet Tuning provides adaptability by means of fine-tuning a sub-network on the encoder in incremental sessions. As shown in Figure 3(a), this trick improves accuracy on novel classes by 11.1%. We also show the accuracy after each session in Figure 3(b). From this figure, we observe comparable accuracies to that of stability tricks in the earlier sessions. However, Incremental SubNet Tuning shows better performance in later sessions when the number of novel classes increases. This again shows the importance of the adaptability trick for learning novel classes. Overall, this trick provides an improved accuracy of 58.12%, 63.1%, and 57.85% on the CIFAR-100, CUB-200, and miniImageNet.
In Figure 3(c), we visualize a t-SNE representation of the learned embeddings (for a randomly selected subset of classes) for the stability tricks and the adaptability tricks. As observed in these figures, stability tricks alone provide sufficient space for the integration of novel classes (5 and 6), although it results in a slight overlap with base classes. In contrast, the incorporation of adaptability tricks creates more separable embeddings.
Training Tricks. Finally, we discuss the results of adding the training tricks. First, we investigate the impact of including the pre-training trick. With the pre-training trick with self-supervised learning, the performance on the three datasets is increased to 58.34%, 63.30%, and 58.01%, respectively. Finally, we observe additional improvements with our last training trick of incorporating the additional learning signal, resulting in accuracies of 58.55%, 63.60%, and 59.11%, on CIFAR-100, CUB-200, and miniImageNet, respectively. To provide an overall evaluation, we present the confusion matrices for both the baseline and the complete bag of tricks in Figure 4. For the baseline, noticeable dark points (correct predictions) are observed along the diagonal for the base classes (first 60 classes), but these become less prominent for the incremental classes. With the bag of tricks, we observe significant improvements for all classes, including novel ones.
4.4 Ablation Study
Table 2: Ablation study.
Stab. Adap. Train. Acc
58.55
58.12
55.85
47.35
51.65
37.34
40.19
43.77
In this section, we present an ablation study on the stability, adaptability and training tricks on the CIFAR-100 dataset. The results of this study are presented in Table 2. It is evident from the table that stability tricks have the highest individual impact on the performance of the model, removing which results in an 11.2% drop in performance. Furthermore, removing stability tricks with any of the other tricks drastically drops the performance, resulting in a final accuracy worse than the baseline. This is caused by the fact that without stability tricks, adaptability tricks alone overfit the new classes, and training tricks alone do not learn generalized features. We find the second important component to be the adaptability tricks, which result in a 2.7% drop in performance when removed.
Table 3: Performance of prior works and our framework for different encoder sizes on CIFAR-100 dataset.
Accuracy
Encoder Baseline Ours
ResNet-20 43.77 58.55
ResNet-50 38.01 58.75
ViT-B/16 32.14 47.89
ViT-B/16 (pre-trained) 50.14 69.72
ViT-B/32 (pre-trained) 50.01 69.58
ViT-L/14 (pre-trained) 51.55 71.09
Performance with Larger Encoders. For all the experiments in this work, we followed the same problem setup and encoders used by all the prior works on FSCIL (Song et al., 2023; Kang et al., 2023). Our focus in this paper is to improve the stability, adaptability, and overall performance of FSCIL while following the same benchmarks and problem setups, as well as encoder backbones, to ensure a fair comparison. However, the tricks in our framework are independent of the encoder’s choice and can be easily adapted to any encoder, including ViT. To investigate the performance of our framework on larger encoders, in Table 3, we show the results on ResNet-50, ViT-B/16, and pre-trained (on ImageNet) ViT-B/16, ViT-B/32, and ViT-L/14. The results from this study show that the proposed tricks not only transfer to larger encoders and ViTs, but the performance gains are even larger than those of smaller networks like ResNet-20. For instance, ResNet-50 shows a 20.74% improvement compared to 14.78% in ResNet-20. Similarly, ViT-B/16 with pre-trained and randomly initialized encoders shows 19.58% and 15.75% improvements over the baseline. Additionally, pre-trained ViT-B/32 and ViT-L/14 show 19.57% and 19.54% improvements over the baseline. This is due to the fact that without the stability and adaptability tricks from our framework, a larger encoder with more parameters is more prone to overfitting, making it difficult to learn incremental classes in a few-shot learning setting. With our proposed framework, the larger encoders show a large improvement in performance.
4.5 Comparison to Prior Works
Finally, we compare the results of our bag of tricks with the prior works on CIFAR-100 in Table 4, and present the results for CUB-200 and miniImageNet in Appendix Tables S1 and S2. The tables present the results for the base session and the overall performance across incremental sessions. The first row in this table presents the results for our baseline. Among prior works, SoftNet held the previous state-of-the-art for CIFAR-100 with a final accuracy of 55.33%. Our framework achieves a performance of 58.55%, marking a 3.22% improvement over the current state-of-the-art. In Figure 5, we plot the accuracy over the incremental sessions for CIFAR-100, CUB-200, and miniImageNet. For CUB-200, the previous state-of-the-art was held by SAVC with an accuracy of 62.50%, which our framework outperforms with an accuracy of 63.60%. Finally, the state-of-the-art on miniImageNet was also held by SAVC, with an accuracy of 57.11%, which our method outperforms by 2.0%, achieving an accuracy of 59.11%.
Method Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
Baseline 70.05 65.37 60.59 56.72 53.38 50.53 48.07 45.87 43.77
Rebalancing (Hou et al., 2019) 61.31 47.80 39.31 31.91 25.68 21.35 18.67 17.24 14.17
iCaRL (Rebuffi et al., 2017) 61.31 46.32 42.94 37.63 30.49 24.00 20.89 18.80 17.21
TOPIC (Tao et al., 2020) 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
IDLVQ-C (Chen & Lee, 2020) 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84
FSLL (Mazumder et al., 2021) 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS (Mazumder et al., 2021) 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M (Shi et al., 2021) 67.28 63.80 60.38 57.06 54.08 51.39 48.82 46.58 44.65
CEC (Zhang et al., 2021) 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
MetaFSCIL (Chi et al., 2022) 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
CLOM (Zou et al., 2022) 74.20 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
C-FSCIL (Hersche et al., 2022) 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
LIMIT (Zhou et al., 2022b) 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
FACT (Zhou et al., 2022a) 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
SAVC (Song et al., 2023) 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
ALICE (Peng et al., 2022a) 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10
SoftNet (Kang et al., 2023) 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
Ours 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.55
Method Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8 9 10
SAVC (Song et al., 2023) 37.74 35.94 34.45 33.16 31.88 30.93 29.80 28.83 28.1 27.37 26.12
SoftNet (Kang et al., 2023) 36.21 34.23 32.35 30.53 28.71 27.15 25.68 24.31 23.16 22.19 21.21
Ours 39.34 37.86 36.49 35.23 34.30 33.83 32.29 31.21 30.77 29.34 28.25
Performance Across Different Shots. Following prior works in FSCIL (Zhou et al., 2022a; Tao et al., 2020), we report the main results in a 5-shot setting. Nonetheless, We also investigate the performance on 1-shot and 2-shot settings to evaluate how the model performs in data-scarce settings and compare the performance with two of the previous state-of-the-art methods, SAVC and SoftNet. We also investigate the 10-shot setting to determine how an increase in data impacts performance. As seen from Figure 6, in 1-shot learning, our proposed framework outperforms prior works by a larger margin than the 5-shot setting. This further shows the effectiveness of the bag of tricks in data-scarce settings. More experiments on 1-, 2-, 5- and 10-shot settings are discussed in Appendix S3.5.
Performance on ImageNet-1K. So far we have report the main results of our approach on, CIFAR-100, CUB-200, and miniImageNet, which are commonly used by prior works in this area. However, these datasets are relatively small in terms of the number of classes. To better understand the scalability of FSCIL in terms of the number of classes, we conduct an experiment on the ImageNet-1K (Russakovsky et al., 2009) dataset, which contains 1000 classes. For this study, we consider 500 randomly selected classes as the base classes and report the results for a 50-way, 10-shot incremental learning setting. The results of this study with our method and its comparison to prior methods are presented in Table 5. As seen from the results, our proposed framework outperforms prior works on the ImageNet-1K dataset by a considerable margin. More specifically, our framework achieves 7.04% and 2.13% improvement over SoftNet and SAVC, respectively.
Table 6: Comparison of computational complexity with a batch-size of 64.
Method Acc. Train Infer. Runtime.
base inc.
SAVC 53.12 810 950 2150 6.1h
Ours (ResNet-18) 58.55 630 935 2150 6.8h
Ours (ResNet-50) 58.75 450 710 1800 8.1h
4.6 Time Complexity
In Table 6, we discuss the time complexity of our framework using a single Nvidia RTX 2080 GPU in comparison to SAVC (Song et al., 2023). To this end, we report the throughput in frames per second (FPS) and the training time. During the training phase on the base session, the throughputs for SAVC and our framework are 630 FPS and 450 FPS, respectively. Nonetheless, during incremental training and inference, our framework with ResNet-18 is as fast as SAVC. Consequently, once trained, our framework can be reliably deployed in real-world applications with the same inference time as previous state-of-the-art while achieving enhanced performance.
Table 7: Performance comparison for base and incremental classes with high semantic similarity.
Method Base Incremental
Baseline 65.3% 19.7%
Ours 74.5% 30.4%
4.7 Discussion
According to the setup adopted in the literature for FSCIL Song et al. (2023); Kang et al. (2023), each of the new classes in the incremental session is different from the classes seen so far. Therefore, incremental classes are generally not exactly identical to the base classes. However, some of the incremental classes can be similar to the base classes. For instance, in CIFAR-100, ‘Orchid’ is a class of flower that is part of the base class, while ‘Sun-flower’ is another flower that is part of the incremental class. A few similar pairs of examples of base classes with similarities to incremental classes are ‘Lion’-‘Tigar’, ‘Bus’-‘Streetcar’, and ‘Shark’-‘Dolphin’, respectively. To understand the effectiveness of the stability on closely related classes, we report the average accuracy on the above-mentioned base and incremental classes with and without our tricks below in Table 7. As we find from this experiment, even for semantically similar incremental classes, our proposed framework is capable of improving the performance of both the base classes (improved stability) and the incremental ones (improved adaptability).
5 Conclusion and Future Work
We present a bag of tricks framework that combines six effective tricks in three distinct categories to improve the stability, adaptability, and overall performance of FSCIL. Stability tricks improve the separation among learned classes to facilitate the learning of new ones, resulting in large improvements on both the base and total classes. Adaptability tricks improve the performance on novel classes by providing more learning capability during incremental sessions. Finally, training tricks provide an additional boost to the final performance. While we did not introduce any new tricks, our contribution in this work lies in developing a framework that combines a set of tricks that jointly improve both stability and adaptability, which, as discussed in the literature review, is a difficult task since improved stability or adaptability often hampers the other aspect. Additionally, We provide extensive analysis of these tricks for a better understanding of their impact beyond the final accuracy, including the impact on stability and adaptability, class separation in the learned embedding space, and performance improvement on base and novel classes. Furthermore, we present a detailed study and comparison to prior works on different aspects that are not explored in the existing literature, including low-shot (1-, 2-shot) performance, results on a dataset with a large number of classes (InageNet-1k), performance on larger encoders (e.g. ResNet-50, ResNet-101), performance on fine-tuning pre-trained foundation models (ViT-B/16, ViT-B/32, and ViT-H/14), performance when new-class has similarity to a base classes (included in revised manuscript). We believe the unified framework and the extensive set of experiments will add value to the further development of the challenging setting of FSCIL.
Though our framework provides adequate adaptability, the performance of the novel classes is still relatively lower compared to the base classes. While this phenomenon has also been widely reported in prior works, our framework was successfully able to reduce the gap between the performance of the base and incremental class from 55% in the literature to 45.2%. However, the performance of the incremental class is still considerably lower than the base class, which can be a potentially interesting direction of inquiry for future work.
Supplementary Material
  • S1: Datasets and Implementation Details
  • S2: Comparison to SOTA
  • S3: Additional Results
    • S3.1: Baseline
    • S3.2: Stability Tricks
    • S3.3: Adaptability Tricks
    • S3.4: Training Tricks
    • S3.5: Performance on Different Shots
Appendix S1 Datasets and Implementation Details
Following previous works (Tao et al., 2020; Song et al., 2023), we evaluate our framework on three popular datasets: CIFAR100 (Krizhevsky et al., 2009), miniImageNet (Russakovsky et al., 2015), and CUB200 (Wah et al., 2011). CIFAR-100 is a dataset of 100 classes, where we use 60 classes (following (Tao et al., 2020)) in the base session and the remaining 40 classes in the incremental sessions. Each incremental session is formulated as a 5-way, 5-shot problem. CUB-200 is a dataset with 200 fine-grained categories. For this dataset, we use 100 classes (following (Tao et al., 2020)) in the base session and the remaining 100 classes in the incremental sessions, with 10 classes in each session. Finally, miniImageNet is a subset of the popular ImageNet dataset that contains 100 classes. For this dataset, we use 60 classes (following (Tao et al., 2020)) for the base session and 5 classes per incremental session over 8 sessions.
For the encoder, we use ResNet-18 by default for all datasets, similar to (Tao et al., 2020; Kang et al., 2023). We train the model with an SGD optimizer, a momentum of 0.9, and a batch size of 64. The learning rate is set to 0.1 for CIFAR-100 and miniImageNet and 0.001 for CUB-200. For all experiments, the model is trained on an Nvidia RTX 2080 GPU.
Appendix S2 Comparison to State-of-the-art
In this section, we present the results for CUB-200 and miniImageNet datasets and their comparison to prior works. As we observe from Table S1, our framework outperforms all existing works on CUB-200 and sets a state-of-the-art of 63.60%. Similarly, for miniImageNet in Table S2, our framework performs better than all prior works, showing a new best accuracy of 59.11%.
@@ 5 Conclusion and Future Work
iCaRL (Rebuffi et al., 2017) 68.68 52.65 48.61 44.16 36.62 29.52 27.83 26.26 24.01 23.89 21.16
EEIL (Castro et al., 2018) 68.68 53.63 47.91 44.20 36.30 27.46 25.93 24.70 23.95 24.13 22.11
TOPIC (Tao et al., 2020) 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28
Rebalancing (Hou et al., 2019) 68.68 57.12 44.21 28.78 26.71 25.66 24.62 21.52 20.12 20.06 19.87
SPPR (Zhu et al., 2021b) 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33
MetaFSCIL(Chi et al., 2022) 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64
F2M (Shi et al., 2021) 81.07 78.16 75.57 72.89 70.86 68.17 67.01 65.26 63.36 61.76 60.26
CEC (Zhang et al., 2021) 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28
FACT (Zhou et al., 2022a) 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94
LIMIT (Zhou et al., 2022b) 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41
SoftNet (Kang et al., 2023) 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75
ALICE (Peng et al., 2022a) 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10
SAVC (Song et al., 2023) 81.85 77.92 74.95 70.21 69.96 67.02 66.16 65.30 63.84 63.15 62.50
Ours 82.31 78.03 75.45 70.99 71.06 67.85 67.44 66.05 64.95 64.31 63.60
@@
iCaRL (Rebuffi et al., 2017) 71.77 61.85 58.12 54.60 51.49 48.47 45.90 44.19 42.71
Rebalancing (Hou et al., 2019) 72.30 66.37 61.00 56.93 53.31 49.93 46.47 44.13 42.19
TOPIC (Tao et al., 2020) 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
EEIL (Castro et al., 2018) 61.31 46.58 44.00 37.29 33.14 27.12 24.10 21.57 19.58
FSLL (Mazumder et al., 2021) 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS (Mazumder et al., 2021) 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M (Shi et al., 2021) 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84
CEC (Zhang et al., 2021) 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63
MetaFSCIL (Chi et al., 2022) 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19
C-FSCIL (Hersche et al., 2022) 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41
FACT (Zhou et al., 2022a) 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49
CLOM (Zou et al., 2022) 73.08 68.09 64.16 60.41 57.41 54.29 51.54 49.37 48.00
LIMIT (Zhou et al., 2022b) 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19
SoftNet (Kang et al., 2023) 79.77 75.08 70.59 66.93 64.00 61.00 57.81 55.81 54.68
ALICE (Peng et al., 2022a) 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70
SAVC (Song et al., 2023) 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11
Ours 84.3 79.59 75.49 71.4 68.45 65.129 62.2 60.52 59.11
Appendix S3 Additional Results
S3.1 Baseline
@@
S3.4 Training Tricks
Finally, we discuss the experiments on the training tricks. In Table S6, we show the results for training different encoders on all the datasets. As we observe in this table, increasing the model size in our framework improves the performance up to a certain model size. Notably, we observe that the best performance for all the datasets is obtained with ResNet-50.
S3.5 Performance on Different Shots
@@
SAVC (Song et al., 2023) 79.85 71.29 64.11 57.08 53.45 50.45 48.74 46.54 44.45
SoftNet (Kang et al., 2023) 79.88 72.01 65.01 59.01 55.01 52.01 50.01 48.01 46.01
@@
SAVC (Song et al., 2023) 79.85 72.00 65.45 59.17 55.25 52.2 50.16 48.11 46.94
SoftNet (Kang et al., 2023) 79.88 73.02 66.01 60.01 56.01 53.01 51.01 49.01 47.01
@@
SAVC (Song et al., 2023) 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
SoftNet (Kang et al., 2023) 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
@@
SAVC (Song et al., 2023) 79.85 75.45 72.18 68.26 65.66 62.15 60.27 58.19 56.01
SoftNet (Kang et al., 2023) 79.88 77.28 73.15 69.34 66.14 63.38 61.27 59.33 57.14